Show that the sound wave intensity I can be written in terms

of the frequency f and displacement amplitude dm in the form, I = 2pi

vf ^2s^2 m.


1
Expert's answer
2021-11-14T17:18:23-0500

Energy of oscillation over time:


"\frac ET=\frac1{2T} mA^2(2\pi f)^2."

Intensity is energy over area:


"I=2\pi^2\rho f^2vs^2."